Loli:

If x^t x^3 = x^6, what is the value of t? A~ 18 B~ 9 C~ 3 D~ 2

Vocaloid:

hint: rule for exponents x^m * x^n = x^(m+n)

Vocaloid:

now let's compare this to your problem x^t * x^3 = x^(6) what do you think t might be equal to?

Loli:

I think 3

Vocaloid:

awesome, 3 is your answer

Loli:

Thanks a bunch!